v1

ATLAS muon chamber construction at NIKHEF

NIKHEF ATLAS MUON GROUP

Contact: Marcel Vreeswijk, h73@nikhef.nl

Abstract: At NIKHEF (Amsterdam), the largest muon chambers for the outer layer of the barrel ATLAS muon spectrometer have been constructed. At this facility, more than 40000 single drift tubes have been produced and 100 drift tube assemblies (muon chambers) have been constructed. We report on the quality assurance of both single drift tube production and chamber assembly. This, together with the also presented results from X-ray scans at CERN (Geneva) and results obtained in a dedicated cosmic ray test set-up, also part of the facility, demonstrates that the chambers fulfil ATLAS specifications.

Contents

1 Introduction 1

2 Coordinate systems 3

3 Drift tubes 3 3.1 Drift tube requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Drift tube production . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Drift tube quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 Drift tube chambers 7 4.1 Assembly station . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2 Chamber assembly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

4.2.1 Results of X-ray scans . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5 Chamber services 13

6 Broken wires 15

7 Tests of completed chambers 15 7.1 Cosmic ray test stand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 7.2 Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

8 Chamber transport to CERN and instalation in ATLAS 17

9 Conclusions 18

A Alignment system RASNIK 21

List of Figures

1 xxxxxxxxxxxxxxxxxxxxxxxxxxxATLAS DETECTOR met MDT ingetekend . . 1 2 xxxxxxxxxxxxxxxxxxxxxxxxxxxMDT principle . . . . . . . . . . . . . . . . . 2 3 Schematic view of a muon chamber. . . . . . . . . . . . . . . . . . . . . . . . . 3 4 Schematic view of a drift tube. Several components are indicated. . . . . . . . . 4 5 Picture of an end-plug and its components. . . . . . . . . . . . . . . . . . . . . 4 6 OVERNIEUW MET ASPECTEN: LEGE HULZEN, EINDPLUGS, DRAAD-

SPOEL, UITEINDELIJKE BUIS. Picture of the wire machine taken at the north side. Indicated are the movable platform in the final position and the clamp to pre-tension the wire. Also indicated is the tube with has fixed end-plugs in this stage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7 Scatterplot of the y and z deviation of the wire position measured on a sub-sample

of produced drift tubes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

8 Leakrate of all used tubes in b l/s . . . . . . . . . . . . . . . . . . . . . . . . . 7

9 a: Dark current in nA of all used tubes. . . . . . . . . . . . . . . . . . . . . . . 8

10 Measurement of the wire tension of all tubes just before chamber assembly. … 9

11 Measurement of the change in wire tension between production of the tube and chamber assembly. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

12 stacking drawing+gluing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

13 OVERNIEW: FOTO of a muon chamber. EVENTUEEL MET DETAILFOTOOS 12

14 Picture taken in the cleanroom, illustrating the last step of the chamber assembly. The jigs on the granite table hold the final tube layer. The chamber, which consistst at this point of the spacer with 3 layers on top and 2 layers at the bottom is lowered on the last tube layer. Also visible are the stacking towers, sag compensation towers and RASNIK towers that control the positioning of the spacer. (xxxxxxxxxx indicate this in pixture? xxxxxxxx) . . . . . . . . . . . . . . . . . 13

15 The picture shows the stacking tower supporting the extension of the spacer. The position of the spacer is monitored the alignment system. xxxxxxxxxindicate in the piture?xxxxxxxx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

16 Tube positions during assembly. . . . . . . . . . . . . . . . . . . . . . . . . . . 15

17 Comparison of results from the tomograph scans for chambers BOL0, BOL2 and BOL3 with RASAS data taken during glueing of the layers of these chambers. For each chamber a set of 4 plots show the deviations from the nominal Z layer pitch for the 3 layers within each multilayer ( MLI and MLII) at the HV and RO side respectively. xxxxxxxxxxxxxxxxxxkeep it? Then, renew itxxxxxxxxxxxxxxxxx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

18 Picture of several alignment components. . . . . . . . . . . . . . . . . . . . . . 17

19 Picture taken the HV side of a BOL chamber. Some individual tubes can be recognized together with the connections to the gasmanifolds. Also shown is a HV board and the backplate of the RASNIK mask component. . . . . . . . . . . 18

20 Picture taken at the RO side of a nearly completed BOL chamber. The RO electronics of the top and bottom multilayer is enclosed in Faraday cages. Also visible are the backplates of the RASNIK cameras, the DCS control box for the temperature and B field sensors, the readout motherboard and two RASMUXes as required for this specific chamber. . . . . . . . . . . . . . . . . . . . . . . . 19

21 Picture of the cosmic ray stand with five BOL chambers. The chamber in the in the highest position is a small type BOL chamber with a width of 48 tubes. … 20

22 Drift time distribution xxxxxxxxxxxxget it, if we want itxxxxxxxxxxxxxxxx . . 20

23 Residual. xxxxxxxxxxxxget plot if we want it.xxxxxxxxxxxxxxx . . . . . . . . 20

24 The principle of RASNIK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

25 The coded mask. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

Photographs to be made:

• Endplug foto with its components • Wiring robot met aspects: lege buis, draadspoel, volle buis, QA • Granieten tafel.

• Granieten tafel detail jigs

• Granieten tafel detail suction

• Granieten tafel detail RASNIK tower

• Granieten tafel detail sphereholders

• Granieten tafel detail spheres etc.

• Granieten tafel guido laser

• Chamber compleet

• Chamber detail faraday cage

• Chamber zonder faraday cage

• Chamber HV cards

• Chamber RO cards=Hedgehogs

• Chamber mezzanine

• Chamber RASMUX

• Chamber TEMPERATURE DCS

• Chamber CSM

• Chamber all sevices

• Chamber gas manifolds

• Chamber gas jumpers

• los hedgehog cars, HV cards, mezzanine,

• aligment stuff

• Chamber als geheel

• cosmic ray stand geheel en details DCS en kamers, trigger Drawings to be made/searched:

• Tube with endplug drawing with twister

• MDT measurement principle

• stacking overview of chamber

• tube gluing of tubes

• granite table with jigs/components (indien al beschikbaar)

• eventueel: stacking steps of chamber

• uitleg sag adjustment

• gas triplet

1 Introduction

Figure 1: xxxxxxxxxxxxxxxxxxxxxxxxxxxATLAS DETECTOR met MDT ingetekend

To stand-alone measure muons that emerge with a transverse momentum of 1 TeV/c with an accuracy \frac{σ_{P_{T}}}{P_{T}} =10% is the most stringent requirement for the muon spectrometer of the ATLAS experiment at the LHC. This translates to 50 μm uncertainty on the sagitta measurement in the 2 Tm integrated magnetic field, provided by an air core toroid.

The muon system in the barrel region (η < 1) is designed such that muon trajectories are measured by three muon chambers as illustrated in figure ??. All three layers consists of an approximately equal number of small and large type chambers. At NIKHEF (Amsterdam, The Netherlands) we constructed 100 ’Barrel Outer Large’ (BOL) chambers, including four spare chambers.

Muons chambers consist of accurately assembled layers of alumnium drift tubes. A drift tube has a diameter close to 3 cm and is closed by two end-plugs that hold a 50 μm thick tungsten wire. Under operational conditions, the tube is filled with an Ar/CO_{2} mixture of 93%/7% and the wire is set a potential of 3xxxxxx V with repect to the grounded tube wall. When a muon traverses the tube, it ionises the gas and the free electrons drift to the wire, generating a signal. The Read-Out electronics measure the arrival time of the signal above an adjustable threshold as shown in figure ??. The arrival time, containing the actual drift time information, is offline used

Figure 2: xxxxxxxxxxxxxxxxxxxxxxxxxxxMDT principle

to determine the drift radius based on a known relation between drift time and radius. Earlier tests show that the drift tubes under well-controlled conditions have single hit resolution of about 80 μm (averaged over drift distance), see [?]. The muon chambers in ATLAS have each at least 2×3 drift tube layers, providing sufficient drift radii to recognize and reconstruct a track segment with high accuracy.

The BOL muon chambers consists of two separate layers (multi-layers), each with three drift tube layers, mounted on an aluminium support structure (spacer) as shown in figure 3. Physical deformations of the chambers during construction, transport and operation are monitored ^{1} with an integrated, so called ’in-plane’, alignment system also indicated in figure 3.

The combination of the uncertainties on single hit resolution, alignment and mechanical precision of the chambers should not exceed 50 μm to meet the aforementioned accuracy on the sagitta measurement. Simulations have indicated that is is required to construct the chambers with a mechanical precision of 20 μm (RMS), which is of unprecented accuracy for such large detectors.

In this report, we describe the production of single tubes and completed BOL chambers, moreover we present the results on the quality assurance operations. Among the tests is the dedicated measurement in the X-ray tomograph at CERN of a limited number (roughly 10%) of the BOL

1For this reason the ATLAS muon chamber are called Monitored Drift Tube (MDT) chambers

Figure 3: Schematic view of a muon chamber.

chambers. Finally, we summarise the muon chamber performance in the cosmic ray set-up at our facility.

2 Coordinate systems

When the chamber is located on the granite table in the ’upward’ position the local chamber coordinates are defined such that the y coordinate points to the sky, the x coordinate runs along the tubes in the horizontal plane. The z coordinate, the precision coordinate, runs perpendicular to the tubes. The x, y and z axis constitute a left-handed coordinate system. The global coordinate system in our cleanroom is then defined by a south-north and west-east direction that runs in the x and z direction respectively. In the ’upward’ position, the Read-out (RO) and High-Voltage (HV) side of the chamber coincides with the south and north respectively. In addition, for engineering purposes, we defined the west side of the setup as the ’reference-side’. The reference side always coincides with the smallest z coordinate of the chamber.

3 Drift tubes

A schematic view of a drift tube is shown in figure 4. A drift tube has a diameter of 29.970 ± 0.015 mm and a wall thichness of 400 μm . The tube is swaged on two precise end-plugs that are manufactured with a precise outer ring of aluminium to position the tube during the assembly in jigs with high precision. The end-plug also connects to the gas system and is closed by a signal cap to connect High Voltage (HV) and Read-Out (RO) electronics boards at the tube’s HV and RO ends respectively.

The 50 μm gold-plated tungsten wire is tensioned at 285 gram and then crimped into the end-

Figure 4: Schematic view of a drift tube. Several components are indicated.

Figure 5: Picture of an end-plug and its components.

plug. A locator (’twister’) with small gas flow resistance, accurately centres the wire with respect to its outer ring.

IETS ZEGGEN OVER BUIS EXTRUSION PROCES en/of STRAIGTNESS VAN BUIZEN????

3.1 Drift tube requirements

The meet the 20 μm precision requirement on the total chamber, the contribution of individual drift tubes has to remain below 10 μm . This imposes the accurate machining of the end-plugs and wire locators. (which determine the position of the wire near the tube ends). Furthermore, to guarantee this precision over the full length of the chamber, the wires need to be tensioned within 285+-15gram with corresponding gravitational sag of 406+-22um. The requirement on the co-centricity of the wire due to non-straightness of the drift tube (anode-cathode distance) is a second order effect and set at 100 μm RMS.

Other than the mechanical precision we impose:

• the dark current of individual drift tubes shall not exceed 10nA at 3300V at 3bar absolute pressure with an Ar/CO_{2} (93%/7%) gas mixture (gas gain 2 × 10^{4}),

• the leak rate under these conditions must be less than 10^{−8} bl/s.

These requirements ensure long term stable operation in the ATLAS experiment.

3.2 Drift tube production

The 40 thousand drift tubes were produced in the temperature-controlled cleanroom parallel to the assembly of the chambers, using a semi-automated wiring machine. On the picture shown in figure 6, taken at the north side (as we defined it), several components are indicated.

Figure 6: OVERNIEUW MET ASPECTEN: LEGE HULZEN, EINDPLUGS, DRAADSPOEL, UITEIN-DELIJKE BUIS. Picture of the wire machine taken at the north side. Indicated are the movable platform in the final position and the clamp to pre-tension the wire. Also indicated is the tube with has fixed end-plugs in this stage.

The machine is loaded by an aluminium tube and two complete endplugs. The wire is transported from its supply by means of airflow through the ’south’ end-plug toward the ’north’ end of the tube. At the north end the wire is literally sucked through the end-plug and fetched by a wire clamp. Now, movable platforms at both sides bring the end-plugs to their final positions in the tube. The two end-plugs are fixed by tube swager using air pressure. At the south end, the wire is fixed at the left end-plug by crimping it and the tube is in the state as shown in the picture. The clamp at the other side moves away to pre-tension the wire by 400 gram before releasing it slowly to its final tension of 285 gram. Crimping the wire at the north end-plug finalises the drift-tube.

3.3 Drift tube quality

The wire tension, length and position was measured immediately after production for every tenth tube to assure proper operation of the wiring machine.

The tension is derived from the measurement of the fundamental harmonic resonance. This measurement is als used in [?] and it is based on the induced response to a sinusoidal current while the tube is partially immersed in a magnetic field.

The wire position with respect to the outer ring of the end-plug was measured using an eletromagnetic method (see [5]). the method uses the response of two coils sensing the field of an sinusoidal current that is driven through the wire. One set of coils is used at each end of the tube and by rotating the tube by 90 degrees two projections (called y and z) for each end are obtained.

Figure 7a shows a scatter plot of these projections. The radial deviation is better than 10 μm

Figure 7: Scatterplot of the y and z deviation of the wire position measured on a sub-sample of produced drift tubes.

(RMS) as appropriate. In these tests, we found only eleven tubes with a deviation larger than 25 μm in one of these projections, which we stored for further (destructive) studies.

The gas tightness and dark current of all produced tubes was measured in a set-up (in the cleanroom) specially designed for this purpose (see appendix ??). This set-up consists of 80 vacumized large tubes (diameter 6xxxx cm) loaded with the drift tubes under test. The drifttubes are filled with 4 bar 1%/99% helium/argon mixture and the wire is set at 3300 V. The leakrate (for Ar/CO_{2} at 3 bar) and dark current measurement is shown in figure 8 and 9 respectively. The results indicate that the majority of the tubes is well within specifications. However, the long tail of the leakrate distribution (not shown) leads to a failure rate of about 1%, making the leakrate the dominant rejection source.

After ageing of at least one month and just prior to using the tubes in the actual chamber assembly, all the tubes were subjected to a second wire tension test. Like the frist measurement, the tension is also derived from the measured resonance frequency, but without the requirement that the tube is immersed in an magnetic field. The electrostatic force due to a supplied sinusoidal HV leads to mechanical wire oscillations. The changing capacity of the tube is measured in an LC circuit, leading to a resonance curve. the method is described in [4]. For our wire with length 4940 mm a weight per unit length of 38.70 mg/m, a resonance frequency of 27.2 ± 0.5 Hz corresponds to 285 ± 15 gram tension. Only 21 tubes failed and the results for tubes that passed are shown in figure 10. As mentioned above, for 10% of the tubes a ’fresh’ measurement of the wire tension is available, allowing to study ’slippage’ that may occur over time. We expect that if any slippage occurs, it occurs dominantly in the in the ageing period and possibly during handling of the tubes. // No evidence for wire slippage was found. The change in resonance frequency exhibits

Figure 8: Leakrate of all used tubes in b l/s

an assymetric shape as shown in figure ??. This suggest a comfortably small wire slippage for a fraction of the tubes.

In total, we have produced about 40000 drift tubes. Our stringent criteria led to a rejection rate of 2.5%. The 100 BOL chambers (96 ATLAS chambers and 4 spares) we produced, contain 38448xxxxxxx tubes that fullfil our requirements.

4 Drift tube chambers

Above we already showed in figure 3 the mechanical layout of a BOL chamber. The aimed mechanical precision of 20 μm RMS is implemented as follows. Any transverse cross section of any chamber can be represented by a single wire grid. The individual wires of a particular chamber are mapped on this grid with residuals distributed within 20 μm RMS. Note that this precision applies to the wires, while the tubes have a looser position requirement of 100 μm RMS (with respect to the wire).

The BOL chambers have a length of 5 m and come in three different widths: 2160 mm (72 tubes/layer), 1680 mm (56 tubes/layer) and 1440 mm (48 tubes/layer), all with two multilayers with 3 tube layers on each side of the spacer. The spacer is constructed of three ’I’ beams, called cross-plates, running orthogonal to the tubes. The cross-plates have low mechanical precision (0.5 mm) and the glue (3M xxxxxxxxx), which connects to the mulilayers, absorbs possible irregularities. The crossplates are themselves connected by two beams, that run longitudinal (which are therefore called long-beams). The horizontal position of the central cross-plate with respect to the long-beams is adjustable to compensate for gravitaional sag of the chamber. For the in-plane alignment system, two camera’s and four masks are mounted on the RO and HV crossplate respectively and the central crossplate houses four lenses. With two parallel and two diagonal light rays several deformations of the spacer and thus chamber can be monitored; most

Figure 9: a: Dark current in nA of all used tubes.

notably the sag and torque of the chamber. The torque is a relative rotation of the two outer crossplates. A total of 10 temperature sensors are mounted on the crossplates to monitor temperature gradients.

Tubes of the same layer lie in the same plane. The tubes of an adjacent layer is shifted by 15 mm (the half of the tube’s width) and stacked as illustrated in figure 12. Tubes of the same layer are grouped in series of three (tripplets) to reduce the number of gas connections to the onchamber manifold. Only the first and last tube in the tripplet connects to the manifold at the RO side and HV side respectively, while the tubes are connected with plasticxxxxx gas jumpers as shown in figure ??. During assembly the length differences of the tubes are absorbed by using jumpers of appropriate thickness (in steps of 100 μm , maximal 300 μm ).

The mounting of electronic boards, equipped with xxxxconnections, is carefully prepared by fixing the rotation of the tubes such that the ground pins finally lie in line within 0.5 mm. In the next chapter we elaborate on the on-chamber services.

Practically, the position requiment of 20 μm on the wires implies that the drift tubes should be accurately assembled into a chamber near the endplugs at the outer cross-plates. The straihgtness of the tubes in between the endplug is of less importance. Therefore, we aimed to set-up an assembly stand based on a granite table with precision mechanics, such as jigs to hold the drift tubes, machined with a precision of 10 μm or better. During assembly, the position of the spacer and individual tubes near the outer cross-plates need to be monitored to the same level. Below we describe the assembly station and the adopted procedure to routinely produce 100 BOL chambers at NIKHEF.

Figure 13 shows a final BOL chamber. Seen...............

Figure 10: Measurement of the wire tension of all tubes just before chamber assembly.

4.1 Assembly station

The gluing of six layers of 72 tubes onto a structure with dimensions 5m × 2m , with a precision of 20 μm is a delicate procedure. Depending on the actual layer to be glued, mistakes are costly in time and material. We constructed an accurate assembly station and adopted a robust procedure to assemble the muon chambers with high precision.

FOTOOS

• chamber

• table, jig , spheres, suction,

• guido, rasnik,

Figure 14 shows the granite table equipped with precision mechanics. The end-plugs are hold by the outer jigs, which are constructed and positioned with a tolerance of 10 μm . The stacking towers have the same accuracy and have adaptable heights to absorb the thickness of the vertical pitch of the tube layers. The central stacking position is shifted by half the horizontal pitch as appropriate. The jigs are fixed at the west side, the reference side, and can expand freely to absorb temperature variations. These components, combined with a stable and uniform temperature (typical variation < 0.5^{∘}C over one day) allow to assemble chambers within the 20 μm RMS precision requirement.

During assembly, extensions with precise spheres (diameter 6cmxxxx) are mounted on the crossplates as can be seen in figure 15. The stacking towers at the reference side support the spheres with so-called sphere-holders that can move freely in the x direction to absorb temperature variations. At the east side, the non-reference side, the sphere-holders can move freely in the horizontal plane. The sphere-holders are designed such that their height during possible horizontal movements is maintained with an accuracy better than a few microns.

Figure 11: Measurement of the change in wire tension between production of the tube and chamber assembly.

When the spacer (with or without already fixed tube layers) is supported by the stacking towers the gravitational of the cross plates is considerably, of the order of 50 μm . To reduce this sag considerably to typically 10 μm the unfinished chamber is carried by an additional system. One of the the eight pneumatic cylinders constituting this system is also visible in figure 15.

4.2 Chamber assembly

FOTOOS

• chamber

• table, jig , spheres, suction,

• guido, rasnik,

The assembly procedure starts with the assembly of the bare spacer, with a ’loose’ accuracy of typically 500 μm . Then, the tubes are attached, layer by layer, using glue. The time for the glue to cure sets the

As already mentioned the quality control during chamber assembly start with re-measuring the wire tension of the tubes to be glued in the particular sequence. After this test, the tubes are positioned in the jigs and the horizontal position of the end-plugs, both at the RO and HV side, is measured using laser beam running parallel to the end-plugs [6]. For this test, a reference rod is placed between two adjacent tubes which partially blocks the laser beam measured by an optical sensor. Deviations in the response respresent a measure of deviations in tube heigts. Note that by measuring between adjacent tubes, the results of this test are somewhat ambiguous case neighbouring tubes are both badly poistioned. In addition, the test is sensitive to tube heights (y)

Figure 12: stacking drawing+gluing

as well as to the distance (z) between tube pairs and the radii of the corresponding end-plugs. On the other hand, a badly positioned tube shows up twice and is tagged easily. In practice, only a few tubes needed to get re-positioned. In some rare cases, we had to replace the complete tube due to, as it turned out, bad swaging of the end-plug. Bad swaging occurred for a particular batch of tubes where the end-plug was not well inserted in the aluminium tube. We aimed at tube height deviations below 20 μm and investigated all tube pairs with responses above 10 μm were inspected by the operators.

The result of these measurements as shown in figure 16, indicate that the tubes are well positioned.

After position of the tubes glue (Araldytexxxxx) is dispensed by means of an automated gluing machine. For the first (and later second layer), glue is dispensed only in between adjacent tubes. The relevant side of the cross-plates is supplied with a sufficient amount of glue (3Mxxxxxx) and the spacer is lowered onto the tube layer. To check the correct positioning of the spacer we used RASNIK systems that monitor the cross-plate extensions at the four stacking towers at the corners of the assembly table. In addition, we used a temporary RASNIK system on both the HV and RO crossplate to monitor their sag. Together with the ’standard’ in-plane alignment components of the spacer, these RASNIK systems provided crucial position information with an precision of 10 μm , which may be regarded as remarkably accurate for such large structure.

PLOTRASNIK, Graziano

After glue curing for at least 12 hours the spacer with the first tube layer is rotated upside down, effectively over its x axis, thus respecting the reference side. The second tube layer is prepared the similar procedure as for the first layer is followed. Now both sides of the spacer have a tube layer attached to it. To glue the third layer, the spacer is lifted and the stacking tower height is adapted to attach the third tube layer. On the tubes of the third layer and onward, two additional glue ropes (Araldytexxxxx) are dispensed to connect adjacent tube layers. To glue the fourth layer, the spacer with its three tube layers is rotated back to its original orientation, limiting the

Figure 13: OVERNIEW: FOTO of a muon chamber. EVENTUEEL MET DETAILFOTOOS

(un)balance of the spacer and tube layers and thus avoids possible effects of asymmetrical weight distribution. After gluing the fifth layer the chamber is turned once more to finish it by attaching the sixth layer.

About every 25-th chamber several checks concerning the accurate positioning of the mechanics were performed. Notably, the distances between the corner stacking towers and the corresponding jig at the RO and HV reference side was measured with an auxiliary device. When these distance deviate, the chamber obtains an unwanted stereo angle build in. Our checks demonstrated that such stereo angle is smaller than 4 μrad over the full period of the production.

4.2.1 Results of X-ray scans

About 10% of all chambers undergo a mechnical precision check in X-ray tomograph [?] at CERN. The tomograph consists of a unit with two stereo X-ray beams that runs over the chamber in the z direction at a fixed x position. Scintillator counters measure the ’shadow-pattern’ of the wires and tube-walls. A dedicated analysis of such pattern results in a wire position map with an accuracy of typical 5 μm . The chamber is considered certified as the distribution of the residuals with respect to the nominal wire map has an RMS below 20 μm . Table 1 lists the RMS values of the residual distribution for the scanned BOL chambers. The numbers indicate that these chamber are assembled according to the specifications.

RMS z μm RMS y μm RMS combined
BOL 2  1000  0000  0
BOL 3  1000  0000  0

Table 1: The table list for each scanned BOL chamber the residuals of the wire coordiantes obtained in the X-ray tomograph. The numbers of the BOL chambers refer to their serial production number.

Figure 14: Picture taken in the cleanroom, illustrating the last step of the chamber assembly. The jigs on the granite table hold the final tube layer. The chamber, which consistst at this point of the spacer with 3 layers on top and 2 layers at the bottom is lowered on the last tube layer. Also visible are the stacking towers, sag compensation towers and RASNIK towers that control the positioning of the spacer. (xxxxxxxxxx indicate this in pixture? xxxxxxxx)

In figure. 17 we compare the layer Z-pitch deviations obtained by the X-ray tomograph with those derived from the RASNIK monitoring at the time the layers were glued. The RASNIK deviations are shifted to set their average deviation to zero as appropriate for this comparison. The distribution of the differences of the two measurements has an RMS of 6 μm , indicating that the quality control during the assembly is well understood.

5 Chamber services

Alignment components As mentioned before, physical deformations of the chambers are monitored with the in-plane alignment system consisting of four RASNIK systems mounted during chamber construction. The inter-chamber alignment consist of RASNIK systems monitoring the position of neighbouring chambers of the same type and RASNIK systems that monitor the projective alignment of the three muon chamber layers. In a test setup at CERN [2] it was demonstrated that an alignment accuracy better than 10 μm (on the change of the relative chamber positions) can be achieved with this system. After chamber construction the mounting platform for these system are glued on the chambers. The components for the in-plane RASNIK systems are shown in figure ??. The power and readout for all alignments systems is provided by an onchamber multiplexer, called RASMUX, which itself is later connected through one D-25 cable to a master system that steers and collects information of several muon chambers. The RASMUX is also shown in figure ??.

Figure 15: The picture shows the stacking tower supporting the extension of the spacer. The position of the spacer is monitored the alignment system. xxxxxxxxxindicate in the piture?xxxxxxxx

Temeprature and magnetic field sensors A total of 16 temperature sensors provide supplementary information on possible chamber deformations. In addition, typical a few magnetic field sensors probe the field for off-line corrections. These sensors connect to the detector control system. The DCS is configured and readout with the CAN protocol.

Gas system After the mechnical completion of the chambers, we started with installing the gas system. The on-chamber gas system consists of two gas manifolds for each multilayer. Tubes of the same layer are grouped in series of three (tripplets) to reduce the number of gas connections to the manifold. Only the first and last tube in the tripplet connects to the manifold at the RO side and HV side respectively. After installing the gas system, the chamber is filled to 3 bar and its pressure (corrected for temprature effects) is monitored for at least 24 hours to assure the gas tightness of the added components. The manifolds have only one input or output and can be quickly connected to the gas service in the experiment.

Electronics All the tubes in the chamber are provided with a ground pin at the RO and HV side respectively. The ends of the ground pins and the pins on the gas caps all lie in the same plane as can be seen in figure ??. The and HV electronic boards are provided with xxxxxxxxxxconnetor to connect to the pins robustly. The digital signal processing and the final time measurement is embedded on so called ’mezzanine’ cards that directly connect to the analogue RO electronics and are housed in individual faraday cages. The digitized signals are shipped to a central on chamber processor card (CSM) as shown in figure 20. The CSM is configured via the DCS and connects to the experiment data acquisistion system by optical fibers.

Figure 16: Tube positions during assembly.

6 Broken wires

WHERE??????????????????????????? After the completion of the first chambers we detected a number of broken wires in several chambers. Despite many dedicated studies, the cause of this problem has not been recovered. However, we observed that broken wires seem to occur only in chambers that have ’seen’ high gas flows (xxxbar/l). Moreover, this problem is restricted to the first (or last) tube of a triplet connecting to the gas manifold. To avoid such high flows we put mechnical flow restrictors in the on chamber gas outlet tubing. In total we have xxxx chambers with broken wires. Usually, the number of broken wires is limited to 1xxxxx, except for one chamber, BOLxxxx, whch has the 24xxxx broken wires. We repair the broken wires, using a procedure that uses the existing o-ring and thus not requires any glue. This comes not for free. The accuracy of the wire position is estimated to a few 100 μm . In principle, the wire position of such wires can be obtained later in the ATLAS experiment with a procedure that uses tracks xxxxxxxxxxxxxxxxmore?.

7 Tests of completed chambers

The main purpose of the test stand is to verify the electronic read-out of each chamber under ATLAS gas (Ar/CO_{2}= 93%/7% at 3 bar) and HV (3300V) conditions.

7.1 Cosmic ray test stand

The cosmic ray test facility at NIKHEF accomodates five BOL chambers and has a fixed trigger system as shown in figure 21. The trigger units (also visible in the figure) consist of two layers of scintillators, separated by 50cm thick iron block to filter out the low energetic muons (i.c.

Figure 17: Comparison of results from the tomograph scans for chambers BOL0, BOL2 and BOL3 with RASAS data taken during glueing of the layers of these chambers. For each chamber a set of 4 plots show the deviations from the nominal Z layer pitch for the 3 layers within each multilayer ( MLI and MLII) at the HV and RO side respectively. xxxxxxxxxxxxxxxxxxkeep it? Then, renew itxxxxxxxxxxxxxxxxx

muon with energies smaller below 0.7 GeV). Since the summer of 2004 till the autum of 2005 we routinely tested BOL chambers; usually we exhanged three chamber middle chamber and kept the outer chamber in place to have a stable reference.

7.2 Results

Figure 22 show a typical difttime distribution for a single wire. The t_{0} and t_{max} for each individual wire is extracted from its drifttime distributions. Large deviations are a good indicator for problems. We often observed a pattern related to groups of three drift tubes constituting a gas-tripplet. After an extended period of flushing (xxxxxxx liter) these structures disappear. One, and only one, severe problem remains, which are the broken wires observed in some chambers as mentioned above. xxxxxxxxxx

Further analysis of the data uses software being developped for ATLAS and involves driftdistance versus drifttime calibration with corresponding track fits to drift circles in individual chambers.

Figure 18: Picture of several alignment components.

The distribution of the residuals for a representative wire is shown in figure 23. The efficiency for single hits is 9xxxx% and the resolution is xxxxxx μm as expected for the current conditions. ONDERWERPEN/FIGUREN:

• hit distributions

• drift properties: TMAX

• efficiencies

• tracks

• problems

• wire/layer locations (correlations XTOMO)

8 Chamber transport to CERN and instalation in ATLAS

ONDERWERPEN:

Figure 19: Picture taken the HV side of a BOL chamber. Some individual tubes can be recognized together with the connections to the gasmanifolds. Also shown is a HV board and the backplate of the RASNIK mask component.

handling tools, common supports, RCP assembly Figures:

• Trasnport frames

• Work at CERN: common support

• BOL above magnet (photo exists already)

• installed BOLs

9 Conclusions

chambers are within specs.

Figure 20: Picture taken at the RO side of a nearly completed BOL chamber. The RO electronics of the top and bottom multilayer is enclosed in Faraday cages. Also visible are the backplates of the RASNIK cameras, the DCS control box for the temperature and B field sensors, the readout motherboard and two RASMUXes as required for this specific chamber.

Figure 21: Picture of the cosmic ray stand with five BOL chambers. The chamber in the in the highest position is a small type BOL chamber with a width of 48 tubes.

Figure 22: Drift time distribution xxxxxxxxxxxxget it, if we want itxxxxxxxxxxxxxxxx Figure 23: Residual. xxxxxxxxxxxxget plot if we want it.xxxxxxxxxxxxxxx

References

[1] The ATLAS MUON TDR, http://atlasinfo.cern.ch/Atlas/GROUPS/MUON/TDR/Web/TDR.html

[2] First System Performance Experience with the ATLAS High Precision Muon Drift Tube Chambers Poster presented by M. Vreeswijk at Vienna WCC 1998. Published in Nucl.Instrum.Meth.A419:336-341,1998 http:atlasinfo.cern.chAtlasGROUPSMUONdatchadatcha.html

[3] M. Woudstra, PhD Thesis, to be published

[4] ELECTROSTATIC DIGITAL METHOD OF WIRE TENSION MEASUREMENT FOR KLOE DRIFT CHAMBER. A. Andryakov et al.. 1998. Prepared for 7th Pisa Meeting on Advanced Detectors: Frontier Detectors for Frontier Physics, La Biodola, Isola d’Elba, Italy, 25-31 May 1997. Published in Nucl.Instrum.Meth.A409:63-64,1998 MDT wire tension measurement using an electrostatic method, ATLAS internal note, ATL-MUON-38-264.

[5] An electromagnetic micrometer to measure wire location inside an ATLAS MDT drift tube, Cambiaghi, M., Cardini, A., Ferrari, R., Gaudio, G., ATLAS Internal Note ATL-MUON-1998-259, 1998 . M. Cambiaghi et al., Proceedings of the 1999 IEEE Nuclear Science Symposium, IEEE Trans. Nucl. Sci. 47 (3) 2000.

[6] ATLAS Internal Note MUON-NO- 29 September 1998 A QA/QC system to monitor the planarity of the tube layers in MDT chambers construction G. Ciapetti, G. De Zorzi, R. Giacalone, L. Montani, D. Pinci Dipartimento di Fisica, Universit di Roma ”La Sapienza” and INFN, Sezione di Roma

[7]

[8] C. Bini et al., Nucl. Instrum. Meth. A 461 (2001) 65.

A Alignment system RASNIK

RASNIK is the acronym for Red ALignment System NIK(HEF). The ’red’ refers to the color of the LEDs used in the very first version invented at NIKHEF. The system consists of a coded mask that is (partially) projected by a lense on a camera as shown in figure 24. The basic binary blocks have dimenesions of 100 μm × 100 μm and form a chessboard like pattern as can be seen in figure 25. An additional code on each nineth row and column, also visible in the figure, allows to reconstruct the coarse alignment. This set-up measures relative transverse movements (X,Y) with respect to the optical axis, the relative rotation (α) over the optical axis of the mask and the camera and the ratio of the distance of the lense-camera and lense-mask (the scale). The resolution depends on several parameters, such as the quality of the lense. Especially for large distances between mask and camera, temperature gradients in the air affect the light path deteriorating the resolution. However, for a setup of 1 m long (in total), a precision of 1 μm on transverse movement, 50 μrad on the rotation and 10^{−4} on the scale is achieved. is measured with a resolution of 50 μrad and the scale to 10^{−}4.

The alignment of the ATLAS muon system is completely based on the RASNIK principle as can be read in detail in ??.

Figure 24: The principle of RASNIK. 22 Figure 25: The coded mask.

23